OpenStudy (anonymous):

please help... The surface area of a right square pyramid whose base has sides of length 12 dm and whose slant height is 11 dm (on all four faces) is?

OpenStudy (anonymous):

the base is a square so Area=side^2 144=12^2 The sides are 2 right angled triangles so Area=base*height 132=11*12 but there are four sides so 132*4=528 plus the base 528+144=672 Answer 672 dm^2

OpenStudy (anonymous):

this really helped

OpenStudy (anonymous):

glad it was helpful